# Non-Iterative Symbolic-Aided Chain-of-Thought for Logical Reasoning

### v1

@@ 6 Conclusion
In this work, we introduced Symbolic-Aided CoT, a novel prompting technique for non-iterative logical reasoning, which achieves superior performance on well-known benchmark datasets—most notably ProofWriter, ProntoQA, and LogicalDeduction. Our method is deliberately simple to preserve generalizability and shows strong potential for extension to other reasoning tasks. For future work, Symbolic-Aided CoT, grounded in structural characteristics, could be combined with mechanisms for refining the latent semantic vector space, thereby further improving the faithfulness and reliability of LLMs’ reasoning capabilities.

### v2

@@ 6 Conclusion
In this work, we introduced Symbolic-Aided CoT, a novel prompting technique for non-interactive logical reasoning, which achieves superior performance on well-known benchmark datasets—most notably ProofWriter, ProntoQA, and LogicalDeduction. Our method is deliberately simple to preserve generalizability and shows strong potential for extension to other reasoning tasks. For future work, Symbolic-Aided CoT, grounded in structural characteristics, could be combined with mechanisms for refining the latent semantic vector space, thereby further improving the faithfulness and reliability of LLMs’ reasoning capabilities.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 22H00524 and the Nakajima Foundation. We used ABCI 3.0 provided by AIST and AIST Solutions with support from “ABCI 3.0 Development Acceleration Use”.
Limitations
We discuss the following limitations and future works: We have evaluated the proposed method on four widely used logical reasoning benchmarks. However, they are mostly synthetic; incorporating real-world datasets or diverse reasoning tasks (e.g., commonsense reasoning) would strengthen claims of generalizability. Relying solely on automatic metrics like accuracy overlooks qualitative aspects; integrating human evaluations to assess reasoning faithfulness and interpretability would offer a more holistic validation. Future research could also explore the method’s robustness to adversarial perturbations, sensitivity to prompts, and scalability to longer reasoning chains.